# VENI, VINDy, VICI – a variational reduced-order modeling framework with uncertainty quantification

### v1

@@ Abstract
The simulation of many complex phenomena in engineering and science requires solving expensive, high-dimensional systems of partial differential equations (PDEs). To circumvent this, reduced-order models (ROMs) have been developed to speed up computations. However, when governing equations are unknown or partially known, or when access to full order solvers is restricted, typically ROMs lack interpretability and reliability of the predicted solutions.
In this work we present a data-driven, non-intrusive framework for building ROMs where the latent variables and dynamics are identified in an interpretable manner and uncertainty is quantified. Starting from a limited amount of high-dimensional, noisy data the proposed framework constructs an efficient ROM by leveraging variational autoencoders for dimensionality reduction along with a newly introduced, variational version of sparse identification of nonlinear dynamics (SINDy), which we refer to as Variational Identification of Nonlinear Dynamics (VINDy).
In detail, the method consists of Variational Encoding of Noisy Inputs (VENI) to identify the distribution of reduced coordinates. Simultaneously, we learn the distribution of the coefficients of a pre-determined set of candidate functions by VINDy. Once trained offline, the identified model can be queried for new parameter instances and/or new initial conditions to compute the corresponding full-time solutions. The probabilistic setup enables uncertainty quantification as the online testing consists of Variational Inference naturally providing Certainty Intervals (VICI). In this work we showcase the effectiveness of the newly proposed VINDy method in identifying interpretable and accurate dynamical system for the Rössler system with different noise intensities and sources. Then the performance of the overall method – named VENI, VINDy, VICI – is tested on PDE benchmarks including structural mechanics and fluid dynamics.
@@
Reduced order modeling , data-driven methods , variational autoencoders , sparse system identification , nonlinear dynamics , generative AI.
@@ 1 Introduction
Scientific computing has emerged as a vital tool across a broad spectrum of applications in engineering and science, facilitating the exploration of complex phenomena that are otherwise intractable. Central to this exploration is the simulation of systems governed by partial differential equations (PDEs), which accurately describe a vast range of physical behaviors. However, a notable challenge arises from the fact that governing equations are not always known or readily available for every phenomenon of interest. Advancements in measurement technology and affordable sensors have enhanced the capability to collect rich spatio-temporal data, thereby fueling significant interest in the automatic extraction of descriptive and predictive physical models directly from data streams. However, even when governing PDEs are known or accurately identified, their computational resolution demands substantial resources. This issue is further compounded in tasks requiring to repetitively solve parametrised PDEs, such as, e.g., in uncertainty quantification (UQ) [1] and shape optimization [2], making high-fidelity simulations computationally impractical for comprehensive analyses or real-time applications.
By combining dimensionality reduction with system in a unified data-driven, non-intrusive framework, we demonstrate a reduced-order model paradigm based on variational inference to perform accurate and uncertainty-aware estimates of full solution fields over time and parameter variations.
Reduced-order models (ROMs) have been developed as a strategic compromise to reduce the computational demand without significantly sacrificing accuracy [3, 4, 5]. Relying on the assumption that most of PDE solution data lie on (or near) a low-dimensional manifold [6], the goal of ROMs is to approximate the high-dimensional solutions through a lower-dimensional representation. In addition to dramatically reducing computational costs, compressing data into a low-dimensional space allows for a non-redundant description of the relevant system [7] and favors the performance of data-driven, system identification techniques [8, 9, 10, 11, 12, 13, 14, 15], as their performance is sensitive to the dimension of the reduced space. This has paved the way for a joint discovery of reduced coordinates and governing equations from data streams, by using, e.g., autoencoders as linear embeddings of nonlinear dynamics [16], or SINDy (Sparse Identification of Nonlinear Dynamics) autoencoders as proposed in Champion et al. [17] with further development to account for partial measurements [18] and for parametric, non-autonomous systems [19]. By explicitly identifying the dynamical model underlying the observed phenomenon, these methods significantly improved the generalization, extrapolation and predictive capability of non-intrusive ROMs, while preserving their flexibility, generability and data-driven nature. However, as these methods rely entirely on data, they inherit the biases present in the dataset. Particularly, their effectiveness may diminish in situations where high-fidelity data are too scarce or costly to obtain to train a sufficiently accurate model. Additionally, when data are marred by significant noise or embody model uncertainties, these methods do not incorporate UQ, compromising the predictive performance and reliability of the estimates.
To address these gaps, we embed the data-driven discovery of latent states and their governing equations in a variational framework with UQ. We present a method which combines variational autoencoders [20] with a newly proposed variational version of SINDy to build a ROM from a limited dataset characterized by high dimensionality and noise. The goal is to exploit the resulting generative ROM to effectively approximate the PDE solution manifold so that we can exploit the generative process to produce reliable solutions.
In detail, the method consists of Variational Encoding of Noisy Inputs (VENI) to identify the distribution of reduced states variables. With respect to their deterministic counterparts, variational autoencoders encourage continuity and smoothness in the learned latent space and promote disentanglement of reduced coordinates [21, 22], demonstrating success in dynamical systems applications [23, 24, 25]. Simultaneously, in a joint offline training, we learn the distribution of the coefficients of a predetermined set of candidate features describing the dynamics by Variational Identification of Nonlinear Dynamics (VINDy). Differently from other approaches which aim at making system identification robust with respect to noise by using Bayesian [26, 27, 28] or weak-form [29, 30] approaches, we employ variational inference which allows for a much more light-weight offline training stage. Variational inference excels as an effective strategy by converting the challenge of UQ into an optimization problem solvable with efficient gradient-based methods. This approach naturally integrates UQ into the training process by adding suitable regularization terms to the loss function, thus preserving the optimization efficiency of neural networks without the need for computationally cumbersome techniques, such as, e.g., Monte Carlo methods. As result of the offline training, we obtain a generative ROM approximating the solution manifold and the latent dynamics. Given a new set of parameters and/or initial conditions, the generative process could be applied to compute full-time solutions and the corresponding prediction uncertainty in an online stage termed as Variational Inference with Certainty Intervals (VICI). A schematic representation of the overall framework – named VENI, VINDy, VICI – is illustrated in Fig. 1
Recent and similar approaches, with slightly different perspective, have been developed to model stochastic dynamics by coupling variational autoencoders with hyper-networks [31], or using stochastic variational inference with Markov Gaussian processes [32]. Other works are pioneering the automatic discovery of latent state variables [7] and governing equations [33] from video data, however relying on non-probabilistic encoding approaches.
@@
The main contributions of the present work are:
@@
    We develop a novel data-driven system identification approach, referred to as VINDy, which is based on variational inference and incorporates uncertainty quantification.
@@
    We integrate dimensionality reduction (VENI), system identification (VINDy), and uncertainty quantification into a unified optimization framework that can be efficiently trained using a variational approach.
@@
    We demonstrate the potential of our approach through various examples, including a low-dimensional chaotic dynamical system and high-dimensional PDE benchmarks, highlighting the method’s ability to identify interpretable dynamics by retrieving the correct governing equations.
This paper is structured as follows. In Section 2 we detail the VENI, VINDy, VICI method in its components. We present how to perform the offline training and how to use the method to efficiently generate full-time solution. Then, we present the potential of the newly introduced VINDy method alone on the Rössler system, thus showcasing the methodology on a didactic, low-dimensional system with different noise intensities and sources (both model and measurement noise). Next we validate the performance of the complete framework through a series of high-dimensional PDE benchmarks, including a straight beam MEMS (Micro Electro-Mechanical System) resonator, which is excited at different forcing amplitudes and frequencies, and an unsteady PDE in fluid dynamics, consisting in a parametrised reaction-diffusion problem. Results for these numerical tests are showcased in Section 3, finally drawing some concluding remarks in Section 4. The source code of the proposed method is made available in the public repository https://github.com/jkneifl/VENI-VINDy-VICI.
@@ 2 Method
2.1 Problem setup
This work presents a method for data-driven system identification and for the efficient generation of solutions of nonlinear, parameterized, time-dependent partial differential equations (PDEs). In particular, we are interested in computing high-fidelity (spatial) approximations of the solutions over a set of $N$ spatial degrees of freedom, which usually come from suitable space discretizations techniques (such as, for instance, finite element, finite volume or isogeometric analysis methods) or represent measurement/sensor locations. The PDE system can then be expressed as a dynamical system of the form
@@ 2.1 Problem setup
where $\bm{x}\in\mathbb{R}^{N}$ is the state of the system, $\dot{\bm{x}}$ its derivative with respect to time $t$ and $\bm{\beta}=[\beta_{1},\ldots,\beta_{p}]^{\top}\in\mathbb{R}^{p}$ the vector collecting $p$ (possibly) time dependent parameters and/or forcing terms. The function $\mathbf{F}$ defines the dynamics of the physical system, evolving from the initial state $\bm{x}_{0}$. As the size $N$ is typically extremely large, system (1) is denoted as full order model (FOM).
Our goal is to compute uncertainty-aware predictions of the time evolution of solutions to (1), given a limited training set of noisy snapshots. We split the objective in two tasks: (i) construct a generative model that allows us to reproduce noise-free, full, spatial states of the system, and (ii) identify the system’s dynamics $\mathbf{F}$ in a probabilistic framework in order to predict the evolution of the states under uncertainty. However, solving these tasks in the high-dimensional space might be extremely complicated and computationally demanding. At the same time, the state dimension $N$ classically results from the numerical discretization scheme or represents measurement degrees of freedom, while often the solution manifold is of significant lower dimension. Consequently, we propose a reduced-order modeling technique that simultaneously reduces the problem’s dimensionality and provides a new set of coordinates better suited to represent the dynamics of the problem, also accounting for uncertainty quantification. More specifically, tasks (i)-(ii) are addressed respectively by the following steps:
  1. (i)
    VENI (Variational Encoding of Noisy Inputs). A generative model based on variational autoencoders (VAEs) is employed to encode the high-dimensional, noisy snapshots data into a more suitable, low-dimensional, latent representation.
  2. (ii)
    VINDy (Variational Identification of Nonlinear Dynamics). A probabilistic dynamical model of the system is learned by a variational version of SINDy (Sparse Identification of Nonlinear Dynamics) [8] on the time series data expressed in the new set of latent coordinates.
Once these two operations are performed offline, we obtain a reduced-order model (ROM) within a probabilistic framework that integrates both the system dynamics and uncertainty quantification. The model can then be queried through an online procedure denoted as
  1. (iii)
    VICI (Variational Inference with Certainty Intervals). Given new parameter/forcing values $\bm{\beta}$ and initial condition $\bm{x}_{0}$, the ROM enables the calculation of the temporal evolution of the system’s solution through variational inference on both the latent variable distribution and the dynamic model. This naturally provides an estimation of prediction reliability through certainty intervals.
The details about the VENI, VINDy, and VICI steps are explained in the following sections.
@@ 2.2 Variational Encoding of Noisy Inputs (VENI)
When relying on data to create models, one cannot assume perfect measurements, but must consider noise that reflects the complexity of the real world, the imperfection of measurements and the lack of knowledge. At the same time, including noise can lead to more robust models and equip them with the ability to cope with unexpected dynamic variations. Consequently, we consider a limited set of noisy snapshot data of a system $\mathbf{X}\coloneqq\{\bm{x}^{(i)}\}_{i=1}^{n_{s}}\in\mathbb{R}^{N}$, which correspond to $n_{s}$ observations distributed according to some unknown distribution $p_{\text{true}}(\bm{x})$ over $\mathbb{R}^{N}$ as starting point for our approach. The strategy of generative models is then to construct a probability distribution $p$, such that $p\approx p_{\text{true}}$. In order to achieve this goal, we make use of variational autoencoders (VAEs), introduced in [20] and comprehensively explained in [34]. VAEs allow variational inference for large datasets and can be used for uncertainty quantification, see [35, 36]. Moreover, they include terms in their objective function that encourage continuity and smoothness in the learned latent space. Accordingly, they allow us to control the way we can model our latent distributions through the choice of suitable priors [33, 26].
The use of autoencoders to create low-dimensional embeddings is justified by the basic assumption of reduced-order modeling: the solutions of PDEs usually lie on a low-dimensional manifold embedded in the high-dimensional, full order space. This allows us to introduce a set of low-dimensional, latent variables $\bm{z}=[z_{1},\ldots,z_{n}]^{\top}\in\mathbb{R}^{n}$, where typically $n\ll N$, from which it is possible to reconstruct the high-fidelity data $\bm{x}$. Specifically, we assume that each high-fidelity data $\bm{x}$ is generated from the latent variables $\bm{z}$ through a decoder mapping $\bm{\phi}_{\text{d}}(\cdot\,;\bm{W}_{\bm{\phi}_{\text{d}}})$, parametrised by a set of weights $\bm{W}_{\bm{\phi}_{\text{d}}}$. The decoder maps latent variables $\bm{z}$ into an output probability distribution over $\mathbb{R}^{N}$:

### v2

# VENI, VINDy, VICI: a generative reduced-order modeling framework with uncertainty quantification

@@ Abstract
Generative models are transforming science and engineering by enabling efficient synthetization and exploration of new scenarios for complex physical phenomena with minimal cost. Although they provide uncertainty-aware predictions to support decision making, they typically lack physical consistency, which is the backbone of computational science. Hence, we propose VENI, VINDy, VICI – a novel physical generative framework that integrates data-driven system identification into a probabilistic modeling approach to construct physically consistent and efficient eeduced-order models with uncertainty quantification.
First, VENI (Variational Encoding of Noisy Inputs) employs variational autoencoders to identify reduced coordinates from high-dimensional, noisy measurements. Simultaneously, VINDy (Variational Identification of Nonlinear Dynamics) extends sparse system identification methods by embedding probabilistic modeling into the discovery process. Last, VICI (Variational Inference with Credibility Intervals) enables efficient generation of full-time solutions and provides uncertainty quantification for unseen parameters and initial conditions. We demonstrate the performance of the framework across chaotic and high-dimensional nonlinear systems.
@@
Reduced-order modeling , data-driven methods , variational autoencoders , sparse system identification , nonlinear dynamics , generative AI.
@@ 1 Introduction
Scientific computing offers a well-estabilished suite of simulation tools that span a wide range of applications in engineering and science, enabling the exploration of complex phenomena that would otherwise be intractable. These phenomena are typically governed by partial differential equations (PDEs), which provide an accurate mathematical framework for their description. However, solving PDEs with high accuracy can be computationally demanding, particularly for high-dimensional systems, complex geometries, or applications requiring repeated evaluations, such as uncertainty quantification (UQ) [1, 2] or PDE-constrained optimization [3].
Reduced-order models (ROMs) have emerged as a promising solution to mitigate these challenges. They represent high-dimensional solutions within a lower-dimensional space, enabling efficient computations, while still retaining the essential accuracy [4, 5, 6]. Traditional ROM techniques rely on detailed knowledge of the governing equations, necessitating domain-specific expertise. This becomes a significant barrier, especially in complex fields like turbulent flow and climate modeling, where the governing equations may be incomplete or unknown. In response, data-driven methods have gained popularity because they can create ROMs directly from observed data, bypassing the need for detailed physical insight. Recent techniques for extracting governing equations from data include dynamic mode decomposition [7, 8] and deep learning approaches based on autoencoders [9, 10, 11, 12, 13, 14], which combine dimensionality reduction with the identification of latent dynamics. However, while these methods improve generalization and predictive capabilities, they often lack robust UQ mechanisms, making them less reliable in noisy or data-limited environments.
Recently, generative models have demonstrated exceptional capabilities in addressing uncertainties and capturing complex patterns in high-dimensional data, e.g. in fields such as computer vision and natural language processing [15, 16]. In the context of ROMs, generative models offer transformative potential by enabling efficient synthesis of new PDE solutions, enhancing robustness to noise, and naturally incorporating UQ. Yet, their application in scientific computing remains limited due to the challenge of ensuring physical consistency and reliability.
To overcome the lack of physical consistency in generative models, we propose a novel physical generative model that seamlessly integrates dimensionality reduction, dynamical system identification, and UQ within a unified variational framework. Our approach embeds latent dynamics discovery directly within the generative model, promoting interpretability and physically meaningful representations of the dynamics using parsimony in sparse regression as a regularizer [17], while retaining the computational efficiency and scalability of variational methods. By combining data-driven identification of reduced variables and governing equations with UQ, this hybrid framework positions our method as a critical advancement for ROMs, enabling automated, interpretable, and uncertainty-aware models that can reliably capture complex system dynamics. The method consists of three steps:
  • (i)
    VENI (Variational Encoding of Noisy Inputs). A variational autoencoder (VAE) [18] is employed to identify the distribution of reduced latent states from high-dimensional, noisy snapshots. VAEs promote continuity and smoothness in the latent space and, when combined with explicit inductive biases, can encourage more independent, disentangled reduced coordinates [19, 20], while not guaranteeing identifiability in the fully unsupervised case [21]. VAEs demonstrated successful applications in dynamical system [22, 23, 24].
  • (ii)
    VINDy (Variational Identification of Nonlinear Dynamics). This novel component extends sparse identification of nonlinear dynamics (SINDy) [25] to a variational framework in order to learn a probabilistic dynamical model in the latent coordinates. Given a library of candidate features, VINDy identifies a sparse subset of active function terms and combines them linearly to approximate the system dynamics. In contrast to SINDy, the multiplicative coefficients are parameterized as probability distributions which capture uncertainty in the dynamics.
VENI and VINDy are trained simultaneously, ensuring that the reduced coordinates learned by VENI enable VINDy to identify a physically meaningful dynamical model. This integrated approach leverages variational inference to achieve efficient and scalable training while avoiding the computational overhead associated with alternative Bayesian [26, 27, 28] or weak-form [29, 30] methods.
  • (iii)
    VICI (Variational Inference with Credibility Intervals). VICI represents the online phase, where the probabilistic ROM generates approximate, full-state solutions and quantifies prediction uncertainty for unseen parameters and/or initial conditions. By leveraging the learned probabilistic representations, VICI ensures reliable extrapolation and provides uncertainty intervals for robust decision-making.
A schematic representation of the overall framework – named VENI, VINDy, VICI – is illustrated in Fig. 1.
@@
Our strategy offers an effective solution by converting the challenge of UQ into an integrated optimization problem solvable with gradient-based methods. This approach naturally encodes UQ into the training process by adding suitable regularization terms to the loss function. This ensures that the optimization efficiency of neural networks is maintained, avoiding the need for computationally intensive techniques like Monte Carlo methods.
Recent work has explored alternative approaches to modeling stochastic dynamics, such as coupling variational autoencoders with hyper-networks [31] or applying stochastic variational inference with Gaussian processes [32]. Additionally, some studies are leading the way in automatically discovering latent state variables [11] and governing equations [27] from video data, although they typically rely on non-probabilistic encoding methods. The present work makes the following contributions:
@@
    We develop a novel data-driven system identification approach (VINDy) which is based on variational inference and incorporates UQ.
@@
    We integrate dimensionality reduction (VENI), system identification (VINDy), and UQ into a unified optimization framework that can be efficiently trained using a variational approach.
@@
    We demonstrate the potential of our approach on the Rössler system, illustrating its performance in a low-dimensional setting with varying noise intensities and sources on a chaotic system, and on high-dimensional PDE benchmarks, including MEMS (Micro Electro-Mechanical System) resonators and an unsteady PDE in fluid dynamics, consisting of a parametrized reaction-diffusion problem.
This paper is structured as follows. In Section 2 we detail the VENI, VINDy, VICI method in its components. We present how to perform the offline training and how to use the method to efficiently generate full-time solution. Then, we present the potential of the newly introduced VINDy method alone on the Rössler system, thus showcasing the methodology on a didactic, low-dimensional system with different noise intensities and sources (both model and measurement uncertainties). Next we validate the performance of the complete framework through a series of high-dimensional PDE benchmarks, including a straight beam MEMS resonator, which is excited at different forcing amplitudes and frequencies, and an unsteady PDE in fluid dynamics, consisting in a parametrised reaction-diffusion problem. Results for these numerical tests are showcased in Section 3, finally we discuss results and draw some concluding remarks in Section 4. The source code of the proposed method is made available in the public repository https://github.com/jkneifl/VENI-VINDy-VICI.
@@ 2 Method
2.1 Problem statement
Our framework is designed to construct a physically constrained generative model for the efficient generation of solutions to nonlinear, parameterized, dynamical systems governed by partial differential equations, given a limited set of noisy snapshots of the system states. Specifically, we aim to compute uncertainty-aware predictions for dynamical systems expressed in the general form
@@ 2.1 Problem statement
where $\bm{x}\in\mathbb{R}^{N}$ is the state of the system, $\dot{\bm{x}}$ its derivative with respect to time $t$ and $\bm{\beta}=[\beta_{1},\ldots,\beta_{p}]^{\top}\in\mathbb{R}^{p}$ the vector collecting $p$ (possibly time dependent) parameters and/or forcing terms. The function $\mathbf{F}$ defines the dynamics of the physical system, evolving from the initial state $\bm{x}_{0}$. As the state dimension $N$ is typically extremely large, system (1) is denoted as full-order model (FOM).
We split the objective into two tasks: (i) construct a generative model that allows us to reproduce full, spatial states $\bm{x}$ of the system, and (ii) identify the system’s governing dynamics $\mathbf{F}$ in a probabilistic framework, enabling predictions of state evolution under uncertainty. However, solving these tasks in the high-dimensional space can be extremely involved and computationally demanding, while often the solution manifold is of a significantly lower dimension [33]. Consequently, our reduced-order modeling technique reduces the problem’s dimensionality and expresses (1) in a new set of reduced coordinates more optimally suited to represent the dynamics of the problem and also accounting for uncertainty quantification. More specifically, task (i)-(ii) are addressed respectively by VENI, which employs VAEs to provide the latent representation and a generative model for the system’s states $\bm{x}$, and VINDy, which identifies the governing dynamics in the new set of latent coordinates.
Finally, in the online phase called VICI, the temporal evolution of the system’s solution is computed for new parameter/forcing values $\bm{\beta}$ and initial conditions $\bm{x}_{0}$, using variational inference on both the latent variable distribution and the dynamic model to ensure robust predictions and reliable uncertainty quantification.
@@ 2.2 Variational Encoding of Noisy Inputs (VENI)
To generate the states $\bm{x}$ of the system, VENI aims to approximate their unknown probability distribution $p(\bm{x})$ over $\mathbb{R}^{N}$, given a limited set of noisy snapshot data of the system $\mathbf{X}\coloneqq\{\bm{x}^{(i)}\}_{i=1}^{n_{s}}$. Given the high dimensionality of the states and the limited amount of available data, this task is particularly challenging. We address it by leveraging variational autoencoders (VAEs), introduced in [18] and comprehensively explained in [34]. VAEs allow variational inference and UQ for high-dimensional datasets  [35, 36], by approximating data distributions in a low-dimensional, latent space. Moreover, using inductive bias in VAEs can promote disentanglement of latent features [21] and allows to control the way latent distributions are modeled through the selection of appropriate priors [27, 26].
Since the solution states $\bm{x}$ of dynamical systems governed by PDEs typically lie on a low-dimensional manifold embedded in the full-order space, we assume that each state $\bm{x}$ can be generated from a set of low-dimensional, latent variables $\bm{z}=[z_{1},\ldots,z_{n}]^{\top}\in\mathbb{R}^{n}$, where $n\ll N$, through the decoder